OpenStudy (flygirlmandi):

find the next three terms in the sequence. 1,1,2,3,5,8,...... A-9,10,11 B-11,14,17 C-13,21,34 D-14,26,50 Thanks :)

OpenStudy (flygirlmandi):

@FaiqRaees

ganeshie8 (ganeshie8):

see any pattern ?

OpenStudy (mud_squad):

Fibonacci series...observe the pattern closely!

OpenStudy (reemii):

^ right. Look that up and compute the next ones.

OpenStudy (reemii):

@flygirlmandi

OpenStudy (mud_squad):

hint- addition

OpenStudy (mud_squad):

ya shuld not have told that mate

OpenStudy (usukidoll):

it does use addition 1 + 1 =2 1+2 = 3 2+ 3 = 5 3+ 5 = 8 so... 8 + 5 = ?????????????????????

OpenStudy (usukidoll):

once you solve 8+5 = ? . add 8 to ?

OpenStudy (flygirlmandi):

guys i am so sorry i fell asleep but i undertand now acording to what u guys said thanks :)